# Multi Task Inverse Reinforcement Learning for Common Sense Reward

### v1

@@ Abstract
One of the challenges in applying reinforcement learning in a complex real-world environment lies in providing the agent with a sufficiently detailed reward function. Any misalignment between the reward and the desired behavior can result in unwanted outcomes. This may lead to issues like “reward hacking” where the agent maximizes rewards by unintended behavior. In this work, we propose to disentangle the reward into two distinct parts. A simple task-specific reward, outlining the particulars of the task at hand, and an unknown common-sense reward, indicating the expected behavior of the agent within the environment. We then explore how this common-sense reward can be learned from expert demonstrations. We first show that inverse reinforcement learning, even when it succeeds in training an agent, does not learn a useful reward function. That is, training a new agent with the learned reward does not impair the desired behaviors. We then demonstrate that this problem can be solved by training simultaneously on multiple tasks. That is, multi-task inverse reinforcement learning can be applied to learn a useful reward function.
@@ 1 Introduction
Reinforcement Learning (RL) is a machine learning paradigm where an agent learns to make decisions by interacting with an environment (Sutton & Barto, 2018). The agent seeks to maximize a cumulative reward signal received in response to its actions. By maximizing this objective, the agent learns a policy that dictates its behavior within the environment. However, in many real-world applications, one needs to design the reward function to precisely define the target behavior. In these scenarios, designing a suitable reward function becomes a key challenge practitioners must address in order to train an agent effectively. Since the agent aims to maximize the reward, any misalignment between the reward and the agent’s intended actions can lead to undesirable behaviors. In the extreme case, this misalignment can lead to “reward hacking” (Skalse et al., 2022) where the agent successfully maximizes the reward without achieving the desired goal. For instance, in (Clark & Amodei, 2022), the authors described a scenario where an agent, trained on the CoastRunners boat racing game learned unwanted behavior. The agent repeatedly knocked out targets in an infinite loop to maximize rewards without ever completing the race. Hence, a crucial element in RL is the design of an effective reward function. This is essential to ensure that agents, trained to maximize this reward, learn the desired behavior, especially when designing agents for operation within complex real-world environments.
@@
To address the reward design problem we first argue that it is natural to split the reward function into two components. One is a simple task-specific reward that defines the goal the agent aims to accomplish. The second is a task-agnostic reward that describes how the agent should behave in the environment while achieving this goal. We refer to this task-agnostic reward as the common-sense reward, cs-reward for short, as it represents “the basic level of practical knowledge and judgment that we all need to help us live in a reasonable and safe way”, using the Cambridge dictionary definition of common sense.
@@
Furthermore, we argue that the tasks-specific reward should be relatively simple to design. For example, consider a scenario where a household robot is assigned chores like throwing garbage and mopping floors. Crafting a reward using computer vision models to verify goal completion is not a significant challenge. However, the cs-reward should be much more complex as it needs to account for a wide variety of cases the agent might encounter. Beyond completing specific tasks, the robot must safely navigate spaces, handle delicate objects with care, conserve electricity, and perform other considerations, such as closing cupboard doors.
@@
Based on this observation, we assume that the reward function can be separated into the task-specific reward, which we assume to be known, and the common-sense reward, which is unknown. We then try to learn the shared common-sense reward from expert demonstrations. A natural approach is to use Inverse Reinforcement Learning (IRL, (Arora & Doshi, 2021) where an agent is trained to imitate the behavior of an expert by simultaneously learning a reward and an agent that tries to maximize said reward. Unfortunately, we show empirically that even when the IRL produces an agent with the desired behavior, it does not learn a meaningful reward. In other words, when we try to train a new agent from scratch using the learned reward, we don’t achieve the desired behavior.
@@
The challenge in deriving a useful reward function, even when IRL successfully imitates the expert, highlights an important distinction between this work and prior works on IRL. Typically, in IRL, the reward serves as a means to an end, i.e., to imitate the expert, and has close connections to the discriminator in Generative Adversarial Networks (GANs) (Ho & Ermon, 2016; Finn et al., 2016). In contrast, this work focuses on the reward itself. Our main question is “will a new agent trained from scratch with our cs-reward gain the designed behavior?”
We aim to address this challenge by leveraging our previous assumptions about the reward structure. Specifically, we utilize multi-task IRL to learn a task-independent shared cs-reward. We term our approach MT-CSIRL. Intuitively, this allows us to combine information from multiple different experts to avoid learning task-specific behavior and spurious correlations. This directs the learned reward to emphasize the underlying shared common-sense reward.
We experiment on the Meta-world benchmark (Yu et al., 2019) using two simple common-sense rewards we designed. We show that even in this simple scenario IRL fails to learn a useful reward and show the importance of multi-task learning over a variety of tasks to learn a useful and transferable reward.
To conclude, one important contribution is the disentanglement of the reward into the task-specific and task-independent components, formulating the common-sense reward. This formulation has many important advantages. First, we can easily combine information from different tasks. Second, it allows us to inject prior knowledge on what each expert is trying to achieve, instead of inferring this from observations as well. Finally, the learned cs-reward can easily be transferred to new unseen tasks. Another contribution is designing a simple MTL approach to learn the cs-reward from expert demonstrations, and evaluating it empirically. Finally, our experiments indicate that training on a wide variety of different tasks is a key component in learning the cs-reward.
Algorithm 1 MT-CSIRL
1:  Input: Expert demonstrations $\mathcal{D}_{t_{1}}...\mathcal{D}_{t_{T}}$, number of iterations $N$, discriminator updates per iteration $D_{\text{updates}}$, Number of tasks $T$, task-specific rewards $r_{t}$, generator updates per iteration $G_{\text{updates}}$
2:  Initialize policies parameters $\omega_{t}$  discriminator parameters $\theta$
3:  for $i\leftarrow 1\textrm{ to }N$ do
4:     for $t\leftarrow 1\textrm{ to }T$ do
5:        for $j\leftarrow 1\textrm{ to }D_{\text{updates}}$ do
6:           Sample trajectories $\tau_{t}\sim\pi_{\omega_{t}}$
7:           Update discriminator $\theta$ using gradient ascent
8:        end for
9:        for $k\leftarrow 1\textrm{ to }G_{\text{updates}}$ do
10:           Sample trajectories $\tau_{t}\sim\pi_{\omega_{t}}$
11:           Update $\omega_{t}$ with $\bar{r}_{t}+\alpha\cdot r_{CS}$
12:        end for
13:     end for
14:  end for
Figure 2: Visualization of train and test tasks for CS-Reward Transfer in MTL CS-Reward transfer experiment
@@ 2 Related Work
Multi-task RL and IRL.
In multi-task learning, a model is trained to perform multiple tasks simultaneously while leveraging shared representations across tasks to improve overall performance and efficiency (Ruder, 2017; Caruana, 1997). Extensive research has been conducted on multi-task RL and IRL, focusing on utilizing shared properties and structures among tasks (Arora et al., 2020; Yang et al., 2020; Vithayathil Varghese & Mahmoud, ; Sodhani et al., 2021; Chen et al., 2023; Zhang et al., 2023), overcoming negative transfer (Yu et al., 2020; Liu et al., 2021; Navon et al., 2022) and performing rapid adaptation to novel tasks (Yu et al., 2021). MTL was also employed in the IRL setting by learning a policy that imitates a mixture of expert demonstrations. MT-IRL generally focuses on the meta-learning setup. Seyed Ghasemipour et al. (2019); Yu et al. (2019) propose learning a policy conditioned on a trained task context vector. Finn et al. (2017b); Yu et al. (2018) propose a MAML-based (Finn et al., 2017a) approach that can adapt a trained policy to a new task with only a few gradient steps. (Xu et al., 2019) proposed to learn a prior over reward functions. In (Rakelly et al., ; Yu et al., 2019) the authors learn a family of rewards by using probabilistic context variables. Differently from these approaches, this paper proposes a method to transfer desired behavior that is not task-specific, without additional adaptation.
Inverse Constrained Learning One line of research that has strong similarities to this work is inverse constrained learning (Malik et al., 2021), and more specifically, multi-task inverse constrained learning (Lindner et al., 2023; Kim et al., 2023). In inverse constrained learning the goal is to learn a set of safety constraints from expert demonstrations. While these safety constraints are conceptually similar to our common-sense reward, our common-sense reward is more general. One can consider a hard constraint as a reward that is $\infty$- ∞ for the forbidden set and zero otherwise. This cannot take into account more subtle effects such as a small negative reward for using up a resource, e.g. energy, that we want to discourage the agent from needlessly wasting, but we do not want to stop it from doing so entirely.
@@ 3 Background
Figure 3: Single task cs-reward: Velocity and action norm during IRL, and during RL training with the learned cs-reward. The red horizontal line represents the target value in the ground truth reward, see Eq. 8 and 9.
@@ 3.1 Inverse Reinforcement Learning
An important RL scenario is imitation learning where the goal is to train an agent on an MDP without having access to the reward, but with expert demonstrations. Commonly, we assume the expert is optimal or near-optimal with regard to the unknown reward. One approach to imitation learning is Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2018), where we imitate the expert by learning a reward and a policy maximizing it to match the expert demonstrations.
In many early IRL approaches, the policy was fully trained at every stage until convergence by using a current reward (Ng & Russell, 2000; Abbeel & Ng, 2004). However, this approach is computationally expensive as it requires training an RL agent from scratch for each reward update. Hence, this approach does not scale well to recent deep reinforcement learning, where the training process can be much more expensive. Current IRL approaches are centered more around training both the reward and the agent simultaneously, updating each one in turn. An important implication of this approach is that it makes the meaning of the learned reward much more uncertain, as there are no guarantees that an agent trained from scratch on the final learned reward will train properly. A visualization of such phenomena is shown in Figure 3. It presents a comparison between training an agent with the learned reward, during IRL process, versus training an agent from scratch using the final learned reward.
(a) Action norm CS reward
(b) Velocity CS reward
Figure 4: Ground Truth CS-Reward: We show the ground-truth common sense behavior on three different experiments. Expert: maximizes task reward and ground-truth cs-reward. MT-CSIRL: trained with the learned cs-reward and with the task reward. SAC: ”vanilla” training, trained only with task reward.
@@
The pioneering works (Ho & Ermon, 2016; Finn et al., 2016) were the first to explore the connection of generative adversarial networks (GANs) to inverse reinforcement learning (IRL), specifically the connection between IRL reward and the GAN discriminator. Their adversarial approach has proven to be the leading method that is been used to train IRL systems (Fu et al., 2018; Jeon et al., 2021; Han et al., 2022). For simplicity, we will base our work on the Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2018) framework, which we found to work well in our experiments.
@@
where $f_{\theta}(s,a,s^{\prime})$ encapsulates the learned reward structure, and $\pi(a\mid s)$ is the policy’s action probability. This formulation leads to a reward function $r_{\theta}(s,a)$ derived as:
@@
4 Our Approach
Our goal is to learn the task-agnostic common-sense reward from expert demonstrations by employing Multi-Task Inverse Reinforcement Learning. We assume each task loss $r_{t}(s,a,s^{\prime})$ has the following structure
@@ 4 Our Approach
and that $\bar{r}_{t}$ is known. This allows for a simple MT-IRL approach where each task policy is trained with a separate weight vector $\pi_{w_{t}}(a|s)$ and the discriminator for each task share weights, capturing the common-sense reward, via
@@
At each iteration, we pick a task $t$, and update the policy based on the following reward:
$r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+f_{\theta}(s,a,s^{\prime}),$ (7)
where $f_{\theta}(\cdot)$ is our learned shared cs-reward. We then update the task-specific discriminator $D^{t}_{\theta}$ with the adversarial loss. Full description of our MT-CSIRL method is provided in Alg. 1.
The key aspect of our approach is the disentanglement between task-specific and task-agnostic rewards. This allows us to train on multiple environments using shared weights for all the different discriminators (see Figure 1). Intuitively, training the shared weights of the different discriminators to simultaneously distinguish expert demonstrations across various environments, avoids learning task-specific behaviors or spurious correlations. We will show empirically in Sec. 5 that training on a variety of tasks is important for learning a transferable common-sense reward function.

### v2

@@ Abstract
One of the challenges in applying reinforcement learning in a complex real-world environment lies in providing the agent with a sufficiently detailed reward function. Any misalignment between the reward and the desired behavior can result in unwanted outcomes. This may lead to issues like “reward hacking” where the agent maximizes rewards by unintended behavior. In this work, we propose to disentangle the reward into two distinct parts. A simple task-specific reward, outlining the particulars of the task at hand, and an unknown common-sense reward, indicating the expected behavior of the agent within the environment. We then explore how this common-sense reward can be learned from expert demonstrations. We first show that inverse reinforcement learning, even when it succeeds in training an agent, does not learn a useful reward function. That is, training a new agent with the learned reward does not impair the desired behaviors. We then demonstrate that this problem can be solved by training simultaneously on multiple tasks. That is, multi-task inverse reinforcement learning can learn a useful reward function.
@@ 1 Introduction
Reinforcement Learning (RL) is a machine learning paradigm where an agent learns to make decisions by interacting with an environment (Sutton and Barto, 2018). The agent seeks to maximize a cumulative reward signal received in response to its actions. By maximizing this objective, the agent learns a policy that dictates its behavior within the environment. However, in many real-world applications, one needs to design the reward function so that it precisely defines the target behavior. In these scenarios, designing a suitable reward function becomes a key challenge that practitioners must address in order to train an agent with the desired behavior. This problem was expressed in Dewey (2014) as the The Reward Engineering Principle: “As reinforcement-learning-based AI systems become more general and autonomous, the design of reward mechanisms that elicit desired behaviours becomes both more important and more difficult.”
Since the agent aims to maximize the reward, any misalignment between the reward and the agent’s intended actions can lead to undesirable outcomes. In extreme cases, this misalignment can lead to “reward hacking” Skalse et al. (2022), where the agent successfully maximizes the reward without achieving the desired goal. For instance, in Clark and Amodei (2022), the authors described a scenario where an agent, trained on the CoastRunners boat racing game learned unwanted behavior. The agent repeatedly knocked out targets in an infinite loop to maximize rewards without ever completing the race. Hence, a crucial element in RL is the design of an effective reward function to ensure that agents, trained to maximize this reward, learn the desired behavior, especially when designing agents for operation within complex real-world environments.
@@
To address the reward design problem, we first argue that there is a natural way to split the reward function into two components. One is a task-specific reward that solely defines the goal that the agent aims to accomplish. The second is a task-agnostic reward that describes how the agent should behave in the environment while achieving this goal. We refer to this task-agnostic reward as the common-sense reward, cs-reward for short, as it represents “the basic level of practical knowledge and judgment that we all need to help us live in a reasonable and safe way”, using the Cambridge dictionary definition of common sense.
@@
Furthermore, we argue that while the overall reward is complex, in many cases the task-specific part should be relatively simple to design. For example, consider a scenario where a household robot is assigned chores like throwing garbage and mopping floors. Crafting a reward using computer vision models to verify goal completion is not a significant challenge. However, the cs-reward should be much more complex as it needs to account for a wide variety of cases the agent might encounter. Beyond completing specific tasks, the robot must safely navigate spaces, handle delicate objects carefully, conserve electricity, and carry out other actions, such as closing cupboard doors.
@@
Based on this, we argue that disentangling the reward is a natural assumption that can aid the reward design problem. Specifically, we propose to separate the reward function into the task-specific reward, which we assume to be known or learned easily, and the common-sense reward, which is unknown. We then try to learn the shared common-sense reward from expert demonstrations. A natural approach is to use Inverse Reinforcement Learning (IRL), Arora and Doshi (2021) where an agent is trained to imitate the behavior of an expert by simultaneously learning a reward and an agent that tries to maximize said reward. Unfortunately, we show empirically that even when the IRL produces an agent with the desired behavior, it does not learn a meaningful reward. In other words, when attempting to train a new agent from scratch using the learned reward, the desired behavior is not achieved.
An important distinction between this work and most prior works on IRL is that we are interested in the reward function itself, while in most cases the reward serves as a tool to imitate the expert. One intuitive explanation as to why IRL fails to learn a useful reward is its strong connections with the discriminator in Generative Adversarial Networks (GANs) Finn et al. (2016); Ho and Ermon (2016). In the ideal case, the GAN discriminator converges to a non-informative constant function. Our main question is “will a new agent trained from scratch with our cs-reward gain the designed behavior?”.
We aim to address this challenge by leveraging our previous assumptions about the reward structure. Specifically, we utilize multi-task IRL to learn a task-independent shared cs-reward. We name our approach MT-CSIRL. Intuitively, this allows us to combine information from multiple different experts to avoid learning task-specific behavior and spurious correlations. This directs the learned reward to emphasize the underlying shared common-sense reward. To show the potential of our proposed disentanglement, we designed two simple synthetic common sense rewards on Meta-world benchmark Yu et al. (2020b). We show that even in this simple scenario IRL fails to learn a useful reward and demonstrates the importance of multi-task learning over various tasks to learn a useful and transferable reward.
To conclude, one important contribution of our work is the proposed disentanglement of the reward into the task-specific and task-independent components, formulating the common-sense reward. This formulation has many important advantages. First, we can easily combine information from different tasks, which we show plays a key role in learning useful reward functions. Second, the learned cs-reward can efficiently be transferred to new tasks. Another important contribution is showing empirically that IRL training might fail to learn a proper reward, despite successfully imitating the expert. We then demonstrate how multi-task learning can play an important role in overcoming this difficulty. Our code is available at: https://github.com/netaglazer/irl-mtl_cs
@@ 2 Related Work
Multi-task RL In multi-task learning, a model is trained to perform multiple tasks simultaneously while leveraging shared representations across tasks to improve overall performance and efficiency Ruder (2017); Caruana (1997). Extensive research has been conducted on multi-task RL and IRL, focusing on utilizing shared properties and structures among tasks Arora et al. (2020); Yang et al. (2020); Vithayathil Varghese and Mahmoud ; Sodhani et al. (2021); Chen et al. (2023); Zhang et al. (2023), overcoming negative transfer Yu et al. (2020a); Liu et al. (2021); Navon et al. (2022) and performing rapid adaptation to novel tasks Yu et al. (2021). MTL was also employed in the IRL setting by learning a policy that imitates a mixture of expert demonstrations. MT-IRL generally focuses on the meta-learning setup. Seyed Ghasemipour et al. (2019); Yu et al. (2019) propose learning a policy conditioned on a trained task context vector. Finn et al. (2017b); Yu et al. (2018) propose a MAML-based Finn et al. (2017a) approach that can adapt a trained policy to a new task with only a few gradient steps. Xu et al. (2019) proposed to learn a prior over reward functions. In Rakelly et al. ; Yu et al. (2019) the authors learn a family of rewards by using probabilistic context variables. Differently from these approaches, this paper proposes a method to transfer desired behavior that is not task-specific, without additional adaptation. Another IRL work that separates between task-specific and task agnostic rewards is Chen et al. (2020). In this work, they propose a method to jointly infer a task goal and humans’ strategic preferences via network distillation. The main difference between the Reward Network Distillation work and our work is, that in their approach they learn the task-specific reward as the shared reward.
Regulated and Constrained IRL One line of research that has strong similarities to this work is inverse constrained learning Malik et al. (2021), and more specifically, multi-task inverse constrained learning Lindner et al. (2023); Kim et al. (2023). In inverse constrained learning the goal is to learn a set of safety constraints from expert demonstrations. While these safety constraints are conceptually similar to our common-sense reward, our common-sense reward is more general. One can consider a hard constraint as a reward that is -$\infty$ for the forbidden set and zero otherwise. This cannot take into account more subtle effects such as a small negative reward for using up a resource, e.g. energy, that we want to discourage the agent from needlessly wasting, but we do not want to stop it from doing so entirely. Another related IRL work is Variational Discriminator Bottleneck Peng et al. (2018). they propose a general technique to constrain information flow in the discriminator by means of an information bottleneck, that can be combined with adversarial inverse reinforcement learning to learn parsimonious reward functions that can be transferred and re-optimized in new settings.
@@ 3.1 Inverse Reinforcement Learning
An important RL scenario is imitation learning where the goal is to train an agent on an MDP without having access to the reward, but with expert demonstrations. Commonly, we assume the expert is optimal or near-optimal with regard to the unknown reward. One approach to imitation learning is Inverse Reinforcement Learning (IRL) Arora and Doshi (2018), where we imitate the expert by learning a reward and a policy maximizing it to match the expert demonstrations. In many early IRL approaches, the policy was fully trained at every stage until convergence by using a current reward Abbeel and Ng (2004); Ng and Russell (2000). However, this approach is computationally expensive as it requires training an RL agent from scratch for each reward update. Hence, this approach does not scale well to modern deep reinforcement learning, where the training process can be much more expensive. Current IRL approaches are centered more around training both the reward and the agent simultaneously, updating each one in turn. An important implication of this approach is that it makes the meaning of the learned reward much more uncertain, as there are no guarantees that an agent trained from scratch on the final learned reward will train properly. A visualization of such phenomena is shown in the Experiment section, in Fig. 2. It presents a comparison between training an agent with the learned reward, during IRL process, versus training an agent from scratch using the final learned reward.
@@
The pioneering works Ho and Ermon (2016); Finn et al. (2016) were the first to explore the connection of generative adversarial networks (GANs) to inverse reinforcement learning (IRL), specifically the connection between IRL reward and the GAN discriminator. Their adversarial approach became the primary approach for training IRL systems Fu et al. (2018); Jeon et al. (2021); Han et al. (2022). For simplicity, we will base our work on the Adversarial Inverse Reinforcement Learning (AIRL) Fu et al. (2018) framework, which we found to work well in our experiments.
@@
where $f_{\theta}(s,a,s^{\prime})$ encapsulates the learned reward structure, and $\pi(a\mid s)$ is the policy’s action probability. This formulation leads to a reward function $r_{\theta}(s,a,s^{\prime})$ derived as:
@@
4 Method
4.1 Preliminaries
Our method relies on the AIRL framework, leveraging its GAN-IRL architecture to infer a reward that maximizes a general common-sense behavior. Differing from AIRL, our method relies on multi-task learning; therefore, we consider a multi-task IRL setup consisting of a set of tasks $\{t;t=1...T\}$, where each task is defined by an MDP $(\mathcal{S},\mathcal{A},\mathcal{T},r_{t},\gamma)$. Each task $t$, has a set of demonstrations $\mathcal{D}_{t}$ from an expert’s policy $\pi^{*}_{t}$. We assume that each expert’s policy maximizes a combination of a task-specific reward, $\bar{r}_{t}$, and a general common sense reward, $r_{cs}$. We aim to recover the task-agnostic common-sense reward, which captures the desired behaviors shared among experts.
4.2 Learning the Common-sense Reward
Our goal is to learn the task-agnostic common-sense reward from expert demonstrations by employing Multi-Task Inverse Reinforcement Learning. Our main assumption is that for each task $t$, the loss $r_{t}(s,a,s^{\prime})$ has the following structure:
@@ 4.2 Learning the Common-sense Reward
and that $\bar{r}_{t}$ is known. This additive split can be modified to other ways to disentangle the rewards, but we use it here for simplicity. This allows for a simple MT-IRL approach where each task policy is trained with a separate weight vector $\pi_{w_{t}}(a|s)$ and the discriminators for each task share weights, capturing the common-sense reward via:
@@
At each iteration, we pick a task $t$, and update the policy based on the following reward $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+f_{\theta}(s,a,s^{\prime})-\log\pi_{t}(a\mid s)$, where $f_{\theta}(\cdot)$ is our learned cs-reward, shared across tasks and designed to capture task-agnostic behaviors.
We then update the task discriminator $D^{t}_{\theta}$ with the adversarial loss, similarly to AIRL:
$\mathcal{L}(\theta)=-\mathbb{E}_{\mathcal{D}_{t}}[\log D^{t}_{\theta}(s,a,s^{\prime})]-\mathbb{E}_{\pi_{\omega_{t}}}[\log(1-D^{t}_{\theta}(s,a,s^{\prime}))],$ (7)
The task discriminators are updated using task demonstrations $\mathcal{D}_{t}$, and trajectories sampled from the current policy $\tau_{t}\sim\pi_{\omega_{t}}$. A full description of our MT-CSIRL method is provided in Alg. 1.
The key aspect of our approach is the disentanglement between task-specific and task-agnostic rewards. This allows us to train on multiple environments while sharing weights among the different discriminators (see Fig. 1). Intuitively, training the shared weights of the different discriminators to simultaneously distinguish expert demonstrations across various environments, avoids learning task-specific behaviors or spurious correlations. We will show empirically in Sec. 5 that training on a variety of tasks is important for learning a transferable common-sense reward function.
Algorithm 1 MT-CSIRL
1:Input: Expert demonstrations $\{\mathcal{D}_{t_{1}},\ldots,\mathcal{D}_{t_{T}}\}$, number of iterations NN, discriminator updates per iteration $D_{\text{updates}}$, number of tasks $T$, task-specific rewards $r_{t}$, generator updates per iteration $G_{\text{updates}}$
2:Initialize: Policy parameters $\omega_{t}$, discriminator parameters $\theta$
3:for $i=1$ to $N$ do
4:  for $t=1$ to $T$ do
5:   for $j=1$ to $D_{\text{updates}}$ do
6:     Sample trajectories $\tau_{t}\sim\pi_{\omega_{t}}$
7:     Update discriminator $\theta$ using gradient ascent
8:   end for
9:   for $k=1$ to $G_{\text{updates}}$ do
10:     Sample trajectories $\tau_{t}\sim\pi_{\omega_{t}}$
11:     Update $\omega_{t}$ using $\bar{r}_{t}+\alpha\cdot r_{\text{CS}}$
12:   end for
13:  end for
14:end for